OpenStudy (anonymous):

The countries with the most cinema screens in the worls are china and the united states. China has 5806 more cinema sceens than the united states, whereas the total screens for both countries is 78,994. Find the number of cinema screens for both countries

OpenStudy (anonymous):

china +US=78994 china-Us=5806 adding them we get 2china=84800 china=84800/2=42400 china+US=78994 42800+US=78994 US=78994-42800=36594

OpenStudy (anonymous):

thank you